# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v3

@@ 2.2.2. Long Context RAG models
Recently, Large Language Model (LLM)-based models have undergone rapid development (Fernandez et al., 2023; Amer-Yahia et al., 2023; Zhang et al., 2024c; Miao et al., 2024; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024b; Ding et al., 2024) and have been successfully adapted to various data mining tasks, including semantic parsing (Drozdov et al., 2022; Mekala et al., [n.d.]; Zhong et al., 2024; Lei et al., 2024; Sinha et al., 2024; Zhou et al., 2024), spatial-temporal mining (Zhang et al., 2024d; Li et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Wang et al., 2024a; Ning and Liu, 2024), and graph mining (Chen et al., 2024b; Tang et al., 2024; Cai et al., 2024; Jiang et al., 2024b; Cheng et al., 2024; Chen et al., 2024a; Fang et al., 2024). Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024), making them originally suitable for dynamic text classification tasks. However, the lack of fine-tuning can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023). To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024). However, the retrieved contents from these models remain unstructured and can be lengthy, which impedes the efficiency and effectiveness of LLMs, leading to lost-in-the-middle issues (Liu et al., 2024a).
2.2.3. Compression-based RAG models
@@
To address the issue suffered by Long Context RAG models, compression-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024b; Wu et al., 2024b; Li et al., 2024a; Zhang et al., 2024a; Jiang et al., 2024a; Guo et al., 2024; Gutiérrez et al., 2024), these models try to compress the side information to reduce the LLM input length. Based on how they compress model inputs, these models can be classified into prompt compressor models, and graph-based RAG models. On the one hand, prompt compressor models (Pan et al., 2024; Jiang et al., 2024a), apply LLM’s generation perplexity on to filter out un-important tokens in the model input. However, these models retrieve side information from each document individually, ignoring the inter-document correlations (Edge et al., 2024; Guo et al., 2024). On the other hand, graph-based RAG models (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024) convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges. Based on the pre-defined retrieval threshold, they then retrieve a limited number of graph nodes and edges to represent the long document. However, the graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks. Also, as the optimal retrieval threshold can vary across different data samples, their globally fixed threshold selected by humans may be suboptimal for the entire dataset.
In this paper, we proposed a novel compression-based graph-RAG model, namely GORAG, which first constructs a weighted graph to model the inter-document correlations and varying confidence and importance of relations, then it achieves adaptive retrieval on this graph, avoiding any human-defined thresholds.
@@ 3.1. Framework Overview
To address the aforementioned uniform-indexing, threshold-dependent, and narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph. As shown in Figure 4, GORAG consists of three core parts, i.e., index graph construction, graph retrieval, and text classification and online Index.
@@
Part 1: Graph Construction. It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round. An example the weighted graph created on WOS dataset is shown in Figure 3, graph formed like this will be used to provide retrieved-augmented information as context for query texts, enabling LLMs to better understand query texts and accurately predict their labels. Specifically, under each new data $\mathcal{D}_{new}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}_{new}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$ from the labeled training texts $\mathcal{T}^{r}_{new}=\{t_{j}\}_{j=1}^{|\mathcal{D}_{new}^{r}|}$. Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}_{new}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}_{new}$. We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance with in the text corpus and relatedness to the label $y$. At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$. More details can refer to Section 3.2.
@@
Part 2: Graph Retrieval. The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all these keywords. From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$. As this MST solely depends on the graph itself and the extracted keywords, we can get rid of any human-defined thresholds and achieve adaptive retrieval. For more details, please refer to Section 3.3.
@@
Part 3: Text Classification and Online Index. In Text Classification part, we first classify the query text with a LLM. This process create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if text formed label names are available. The LLM will carry out classification with this input. After classification, the online indexing procedure dynamically indexes keywords, denoted as $\mathcal{V}_{\text{notexist}}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$. We then integrate these keywords into the graph to further enrich its structure, with edge weights assigned based on their importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$. It enhances the model’s ability to make more accurate predictions. For more details, please refer to Section 3.4.
3.2. Part 1: Graph Construction
@@
In this subsection, we introduce the graph construction procedure of GORAG, which includes a novel edge-weighting mechanism to address the uniform indexing issue. The pseudo code of graph construction is shown in Algorithm 1 in Appendix Appendix C.
For each round $r$, given its respective labeled training texts $\mathcal{D}_{new}^{r}=\cup_{y_{i}\in\mathcal{Y}_{\text{new}}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}_{new}$ from $\mathcal{D}_{new}^{r}$. Specifically, it uses the LLM model with an extraction instruction prompt $p_{extract}=$Please extract some keywords from the following passage” for this extaction. Then, we can get all text keywords $\mathcal{K}^{r}_{new}$ based on texts $\mathcal{D}^{r}_{new}$ as follows.
(2) $\mathcal{K}^{r}_{new}=\bigcup_{t\in\mathcal{D}^{r}_{new}}LLM(p_{extract},t).\vspace{-5px}$
These keywords are served as graph nodes, we denote them as keyword nodes.
Also, GORAG incorporats new candidate labels $\mathcal{Y}^{r}_{new}$ at the $r$-th round as graph nodes, denoted as label nodes. Hence, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
(3) $\mathcal{V}_{new}^{r}=\mathcal{K}_{new}^{r}\cup\mathcal{Y}_{new}^{r}.$
Then, GORAG add edges between each keyword node $v\in\mathcal{K}_{{new}}^{r}$ to its corresponding label node $y\in\mathcal{Y}_{{new}}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
Next, considering the keywords $\mathcal{K}_{new}^{r}$ from the labeled text are not uniformly related to the text’s label $\mathcal{Y}_{new}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge. This weight can be regarded as the correlation between keywords and each label. Firstly, we apply the normalized TF-IDF score (Sammut and Webb, 2010) as our correlation score $CS(v,t)$ to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}_{new}^{r}$ w.r.t. the respective text $t$ it extracted from:
(4) $\vspace{-5px}CS(v,t)=\frac{count(v,t)}{|t|}\times log\frac{|\mathcal{T}^{r}|}{1+|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|},$
@@
where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}_{new}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$.
@@
(5) $\vspace{-3px}w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,y,v}}1-CS(v,t_{j})}{|\mathcal{T}^{r,y,v}|},$
where $\mathcal{T}^{r,y,v}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$. We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively, an example of the generated graph is shown in Figure 3.
Lastly, $\mathcal{G}^{r}_{new}$ will be merged into the graph from previous rounds $\mathcal{G}^{r-1}$ to form the full graph $\mathcal{G}^{r}$ at round $r$:
@@
Particularly, we define $\mathcal{G}^{0}=\emptyset$ and $\ r\geq 1$. Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$ and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
(7) $\vspace{-5px}\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vspace{-5px}$
(8) $\vspace{-3px}w_{y_{n},y_{o}}^{r}=\frac{1}{2}\big{(}\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\big{)}$

### v4

@@ 2.2.2. Long Context RAG models
Recently, Large Language Model (LLM)-based models have undergone rapid development (Zhang et al., 2024b; Snyder et al., 2024; Zhong et al., 2024; Yan et al., 2024; Chen et al., 2024a; Ding et al., 2024). Notably, LLMs are inherently capable of inference without fine-tuning (Sun et al., 2021; Jiang et al., 2023; Touvron et al., 2023; Achiam et al., 2023; Yang et al., 2024a), making them originally suitable for few-shot tasks. However, the lack of fine-tuning can lead LLMs to generate unfaithful responses, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations (Zhang et al., 2023). In web for good applications such as hate speech detection, hallucinations can lead to failure to detect harmful responses, making these harmful contents being spread through internet. To mitigate hallucinations, researchers try to retrieve text documents outside of LLM as side information to help the LLM inference, namely Long Context RAG models (Chen et al., 2023; Sarthi et al., 2024; Hu et al., 2024; Wu et al., 2024a; Cai et al., 2024; Oosterhuis et al., 2024; Yadav et al., 2024; Che et al., 2024; Pan et al., 2024; Jiang et al., 2024). Among them, prompt compressor models (Pan et al., 2024; Jiang et al., 2024), apply LLM’s generation perplexity on to filter out un-important tokens in the model input.
2.2.3. Graph-based RAG models
@@
However, the retrieved contents of Long Context RAG models remain unstructured and lengthy, making the important fine-grained entity-level information be overwhelmed by noisy contexts (Edge et al., 2024; Guo et al., 2025), which still impedes the effectiveness or efficiency of LLM classification, leading to lost-in-the-middle issues (Liu et al., 2024b) for cheaper open-sourced LLMs, or high token cost for proprietary LLMs (Hurst et al., 2024). To address these issues, Graph-based RAG models were proposed (Peng et al., 2024; Zhang et al., 2024a; Wu et al., 2024b; Zhang et al., 2024c; Guo et al., 2025; Jimenez Gutierrez et al., 2024; gutiérrez2025ragmemory), these models convert text chunks from the side information document as graph nodes, and the relation between these text chunks as graph edges. These models then retrieve graph nodes and edges as the LLM classification context, better preserving more fine-grained, entity-level information. However, current Graph-based RAG models usually consider each graph edge uniformly and ignore the varying confidence and importance of relations between text chunks. Also, as the optimal retrieval threshold can vary across different data samples, their globally fixed threshold selected by humans may be suboptimal for the entire dataset.
@@ 3.1. Framework Overview
To address the aforementioned Uniform-indexing, Non-adaptive Retrieval, and Narrow-source issues, we propose GORAG, a novel approach that achieves adaptive retrieval by extracting valuable side information from a minimum-cost spanning tree generated on the constructed graph. As shown in Figure 2, GORAG consists of three core parts, i.e., Offline Graph Indexing, Graph Retrieval, and Classification & Online Graph Indexing.
@@
Part 1: Offline Graph Indexing. It targets to construct the weighted graph $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ based on the labeled data $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$ at the $r$-th round. An example weighted graph created is shown in Figure 2, this graph will be used to provide retrieved-augmented information as contexts for query texts, enabling LLMs to better understand query texts and accurately classify them. Specifically, with the new labeled data $\mathcal{D}^{r}$ at the $r$-th round, we first extract keywords $\mathcal{K}^{r}\in\{t_{j}\}_{j=1}^{|\mathcal{D}^{r}|}$. Then, we will assign an edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ between each keyword $v\in\mathcal{K}^{r}$ and their respective text’s label $y\in\mathcal{Y}^{r}$, where $\mathcal{Y}^{r}=\bigcup_{i=1}^{r}\mathcal{Y}^{i}_{new}$ denotes the union label set of all new labels $\mathcal{Y}^{i}_{new}$ from round $i\in\{1,2,\dots,r\}$. We then compute the weight $w^{r}_{v,y}\in\mathcal{W}^{r}_{new}$ for each edge $e^{r}_{v,y}\in\mathcal{E}^{r}_{new}$ based on the keyword’s importance within the text corpus and keyword relatedness to the label $y$. At last, we merge all graphs $\mathcal{G}^{r}_{new}(\mathcal{V}^{r}_{new},\mathcal{E}^{r}_{new},\mathcal{W}^{r}_{new})$ at each round $r$ as the full graph $\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{W}^{r})$. More details can refer to Section 3.2.
@@
Part 2: Graph Retrieval. The Graph Retrieval process maps the extracted keyword nodes $\mathcal{V}^{t}$ from the query text $t$ to the constructed weighted graph $\mathcal{G}^{r}$ and generates a minimum-cost spanning tree on $\mathcal{G}^{r}$ that includes all mapped keywords in $\mathcal{V}^{t}$. From this minimum-cost spanning tree (MST), the candidate label set $\hat{\mathcal{Y}}_{t}^{r}$ at the current round $r$ is obtained, which is a reduced subset of the original target label set $\mathcal{Y}^{r}$ at the current round $r$. As this MST solely depends on the graph itself and the text extracted keywords, we can eliminate any human-defined thresholds and achieve adaptive retrieval. For more details, please refer to Section 3.3.
@@
Part 3: Classification & Online Graph Indexing. In the Classification part, we create LLM input with the query text $t$, the candidate labels $\hat{\mathcal{Y}}_{t}^{r}$ retrieved from the weighted graph $\mathcal{G}^{r}$, and descriptions $\mathcal{K}_{y_{i}}$ associated with each candidate label $y_{i}\in\hat{\mathcal{Y}}_{t}^{r}$ if available. The LLM will carry out classification with this input. After classification, the online graph indexing procedure dynamically adds keywords $\mathcal{V}_{notexist}^{t}$, which are extracted from the query text $t$ but are not in the existing graph $\mathcal{G}^{r}$, as new graph nodes to further enrich the graph. We then assign edge weights based on these newly added keywords’ importance within the text corpus and their relatedness to the predicted label $y^{*}_{t}$ for the query text $t$. By penalizing the edges between nodes and their semantically irrelevant labels, GORAG can reduce the error introduced by inaccurate classification. The Online Graph Indexing procedure improves the model classification ability by updating the graph in real-time with new keywords. For more details, please refer to Section 3.4. Due to space limitations, we put the time and space complexity analysis of GORAG to Appendix D.
3.2. Part 1: Offline Graph Indexing
@@
In this subsection, we introduce the graph construction procedure of GORAG, which includes a novel edge-weighting mechanism to address the uniform indexing issue. The pseudo code of graph construction is shown in Algorithm 1 in Appendix C.
For each round $r$, given its respective labeled training texts $\mathcal{D}^{r}=\cup_{y_{i}\in\mathcal{Y}^{r}}\{t_{j},y_{i}\}_{j=1}^{k}$, GORAG first extracts text keywords $\mathcal{K}^{r}$ from $\mathcal{D}^{r}$. Specifically, it calculates the normalized TFIDF score (69) as the correlation score $CS(v,t)$ for texts’ each word and phrase with less than 3 words (Wang et al., 2025).
(2) $\vskip-4.015ptCS(v,t)=\frac{count(v,t)}{|t|log(|\mathcal{T}^{r}|+1)}\times log\frac{|\mathcal{T}^{r}|+1}{|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|+1},$
@@
where $\mathcal{T}^{r}=\cup_{i=1}^{r}\mathcal{T}^{i}$ denotes all training and query texts seen so far, $count(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}^{r}|$ denotes the number of texts in the corpus $\mathcal{T}^{r}$ that contain the keyword $v$. Here, we will extract words or phrases with $CS(v,t)>\tau$ as our keyword. We set $\tau=0.5$ following the keyword extraction setting in (Wang et al., 2025). These keywords serve as graph nodes, and we denote them as keyword nodes.
Also, GORAG incorporates new labels $\mathcal{Y}^{r}_{new}=\mathcal{Y}^{r}\setminus\mathcal{Y}^{r-1}$ at the $r$-th round as graph nodes, denoted as label nodes, where $\mathcal{Y}^{r}$ is the total label set at round $r$. Hence, the graph node set $\mathcal{V}^{r}_{new}$ at the $r$-th round can be obtained as follows.
(3) $\vskip-5.01874pt\mathcal{V}_{new}^{r}=\mathcal{K}^{r}\cup\mathcal{Y}^{r}_{new}.$
Then, GORAG adds edges between each keyword node $v\in\mathcal{K}^{r}$ to its corresponding label node $y\in\mathcal{Y}^{r}$, indicating that the keyword $v$ appears in texts associated with the label $y$.
Next, considering the keywords $\mathcal{K}^{r}$ from the labeled text are not uniformly related to the text’s label in $\mathcal{Y}^{r}$, we apply an edge weighting mechanism to assign a weight $w_{v,y}$ to each keyword-label edge between any two nodes $v$ and $y$. This weight can reflect the correlation between keywords and each label. Firstly, we apply the correlation score $CS(v,t)$ assigned by Equation (2) to measure the importance and relatedness of a particular keyword $v\in\mathcal{K}^{r}$ w.r.t. the text $t$ it extracted from.
@@
(4) $w_{v,y}^{r}=\frac{\sum_{t_{j}\in\mathcal{T}^{r,v,y}}1-CS(v,t_{j})}{|\mathcal{T}^{r,v,y}|},$
where $\mathcal{T}^{r,v,y}=\{t_{j}|v\in t_{j}\land t_{j}\in\mathcal{T}^{r,y}\}$ is the text that contains keyword $v$ and labeled $y$. We denote the generated weighted graph for $\mathcal{K}_{new}^{r}$ and $\mathcal{Y}_{new}^{r}$ at the $r$-th round as $\mathcal{G}^{r}_{new}(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r},\mathcal{W}_{new}^{r})$, where $\mathcal{V}^{r}_{new}$, $\mathcal{E}^{r}_{new}$, and $\mathcal{W}^{r}_{new}$ denotes the node, edge and edge weight set respectively. Lastly, $\mathcal{G}^{r}_{new}$ will be merged into the graph from previous rounds $\mathcal{G}^{r-1}$ to form the full graph $\mathcal{G}^{r}$ at round $r$:
@@
Particularly, we define $\mathcal{G}^{0}=\emptyset$ and $r\geq 1$. Also, to guarantee the graph connectivity of the resulting graph, we add edges between every new label node $y_{n}\in\mathcal{Y}_{new}^{r}$, and each old label node $y_{o}\in\mathcal{Y}^{r-1}$, the edge weight $w_{y_{n},y_{o}}^{r}$ of edge $e_{y_{n},y_{o}}$ at round $r$ is defined as the weighted average of all edge weights that link keyword nodes with label node $y_{n}$ or $y_{o}$:
(6) $\vskip-5.01874pt\mathcal{M}_{y}^{r}=\{v\ |\ v\in\mathcal{N}^{r}(y)\ \land\ v\notin\mathcal{Y}^{r}\}\vskip-5.01874pt$
(7) $\vskip-3.01125ptw_{y_{n},y_{o}}^{r}=\frac{1}{2}\Big(\frac{\sum_{v\in\mathcal{M}_{y_{n}}^{r}}w_{v,y_{n}}^{r}}{2\times|\mathcal{M}_{y_{n}}^{r}|}+\frac{\sum_{v\in\mathcal{M}_{y_{o}}^{r}}w_{v,y_{o}}^{r}}{2\times|\mathcal{M}_{y_{o}}^{r}|}\Big),$